500 liters of gas in a flexible-walled container are prepared at 700 mmHg and 200°C. The gas is placed into a tank under high pressure. When the tank cools to 20°C, the pressure of the gas is 30 atm. What is the volume of the gas?

1
Expert's answer
2022-03-13T15:20:21-0400

Temperatures in Kelvins:

200oC = 273 + 200 = 473 K

20oC = 273 + 20 = 293 K


700 mmHg = "\frac{700}{760}=0.921\ atm"


According to the combined gas law,


"\frac{P_1V_1}{T_1}=\frac{P_2V_2}{T_2}"


Therefore,


"V_2=\frac{P_1V_1T_2}{T_1P_2}=\frac{0.921\ atm\times500\ L\times293\ K}{473\ K\times30\ atm}=9.5\ L"


Answer: 9.5 L